Ann gets a 30 year 1/1 Fully Amortizing ARM for $1,000,000, with monthly payments and monthly compounding. The initial rate is 3%. In the future, the rate will reset to 250 basis points above the LIBOR. There are no rate caps or floors. Suppose at the first reset, the LIBOR was 6%. What is the monthly mortgage payment for the second year?"
Answer #1

USING FINANCIAL CALCULATOR
Step 1: Monthly payment initially
N=12*30
I/Y=3%/12
PV=-1000000
FV=0
CPT PMT=4216.0403372945

Step 2: Loan balance after 1st year
N=12*1
I/Y=3%/12
PV=-1000000
PMT=4216.0403372945
CPT FV=979121.996415604

Step 3: New monthly payment
N=12*29
PV=-979121.996415604
I/Y=(6%+2.5%)/12
FV=0
CPT PMT=7585.94797296199

USING EXCEL
=-PMT((6%+2.5%)/12,12*29,FV(3%/12,12*1,PMT(3%/12,12*30,-1000000),-1000000))
=7585.95